Opened 5 years ago

Closed 4 years ago

#524 closed flight processing (fixed)

RG12/09, flight day 159/2014, Wessex

Reported by: knpa Owned by:
Priority: alpha 5 Milestone: 2014 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by tec)

Data location: /users/rsg/arsf/arsf_data/2014/flight_data/uk/RG12_09-2014_159_Wessex/

Data arrived from ARSF via network transfer on 09/06/2014
Data archived on 13/01/2014

RG12/09 Scientific objective: Biodiversity and the provision of multiple ecosystem services in current and future lowland multifunctional landscapes

RG12/09 Priority: 10

PI: France Gerard

Sensors:

Camera (Requested), delivered 14/08/2014
Fenix (Requested)
LiDAR (Requested), delivered 14/08/2014
FW LiDAR
OWL

Change History (30)

comment:1 Changed 5 years ago by tipo

Started navigation processing.

comment:2 Changed 5 years ago by tipo

Finished navigation, started hyperspectral processing.

comment:3 Changed 5 years ago by dap

Started RCD processing.

comment:4 Changed 5 years ago by dap

Started LiDAR processing.

comment:5 Changed 5 years ago by dap

Creating RCD delivery.

comment:6 Changed 5 years ago by dap

RCD delivery creation finished. Awaiting delivery check.

comment:7 Changed 5 years ago by dap

A roll error of -0.005892596 seems correct for lines 085223 and 090210. Now checking for consistency across all lines.

comment:8 Changed 5 years ago by stgo

Picking up hyperspectral from tipo, he's produced the SCT values already. I will reprocess the lines to update to the new sct values from apl.

comment:9 Changed 5 years ago by dap

Inconsistency found on other lines with previously used roll error. Looking for roll error using lines 090210 and 091133

comment:10 Changed 5 years ago by mark1

  • Description modified (diff)

comment:11 Changed 5 years ago by dap

Consistent roll error of -0.005732596 found for all lines. Now classifying LAS files.

comment:12 Changed 5 years ago by mark1

RCD delivery check complete.

Re-created the KML file as some of the file paths and line naming were a bit odd. Otherwise OK.

Ready to deliver.

comment:13 Changed 5 years ago by dap

Finished manually classifying LAS files and copying classifications to full waveform data. Now making lidar delivery.

comment:14 Changed 5 years ago by dap

Moved all full waveform related files to processing/tmp_fw/ temporarily to create delivery. Will move files back when discrete delivery is created.

comment:15 Changed 5 years ago by dap

Moved all full waveform data back to its proper directory from the temporary directory, per the file name conventions.

LiDAR processing (incl. full waveform) complete, awaiting delivery check.

comment:16 Changed 5 years ago by dac

Starting delivery check of LiDAR data.

comment:17 Changed 5 years ago by dac

Delivery check OK. Some minor issues identified and fixed. Missing screenshot with vectors - check if required.

Fixed:

  • Extra tif in screenshots file - deleted
  • full_waveform_lidar.pdf missing from docs - copied from /users/rsg/arsf/doc/data_quality_reports/
  • Problem with LAS 1.3 link and LaTeX double quotes in Readme file - have corrected and fixed generation script.
  • ASCII LiDAR data didn't have CRLF line terminators - fixed with unix2dos
  • sol file incorrectly named - have corrected.

Other

  • Lots of cloud cover (20 - 30 %). The majority of points had been classified as noise, spotted a very small number of unclassified points, had been noted in read me

comment:18 Changed 5 years ago by dap

Spoken to mark1 about vectors for this project, PI does not require vectors. LiDAR ready for delivery.

comment:19 Changed 5 years ago by dap

LiDAR (including full waveform) and camera data dispatched to France Gerard on 14/08/2014.

comment:20 Changed 5 years ago by dap

  • Description modified (diff)

comment:21 Changed 5 years ago by stgo

Hyperspectral data ready for delivery check, note these lines have been split.

comment:22 Changed 5 years ago by tec

Fenix Delivery Check
Starting delivery check.

comment:23 Changed 5 years ago by tec

Fenix Delivery Check
Two hyperspectral deliveries. One needs removing, checking 'RG12_09-159-hyperspectral-20140827'

Readme flightline name to fenix file id table is split over two pages, would it be better to move it so its all on the next page?

Logsheet page 1 is messed, the length of the data quality field is moving the right edge of the table off the page. Logsheet does not contain any notes for the flightlines or general notes.

In 'flightlines' empty fodis directory/

No mosaic with vectors

8 Flightlines but 22 screenshots, assuming some lines have been split. All 'flightlines' sub directories contain multiples of 22.

'fenix_viewvector_list.txt' does not use windows line terminators, run

unix2dos sensor_FOV_vectors/fenix_viewvector_list.txt

All screenshots have blue spots on them, mainly where there is cloud.
Mosaic image should be called fenix_mosaic.jpg not mosaic_image.jpg, needs to be recreated anyway.

Level1b files:
f159131b, f159141b >90% underflows around band 500
f159151b, f159221b >50% underflows around band 500

Level 3 files have NOT been zipped yet. Will wait until problems have been corrected.

Delivery Check finished for now.

comment:24 Changed 5 years ago by tec

Fenix
Fixed fenix_viewvector_list.txt's line terminators, they are now CRLF.

Logsheet is also fixed.

comment:25 Changed 5 years ago by tec

Fenix
Zipping the mapped files.

comment:26 Changed 5 years ago by tec

Fenix
Sent off to France at 14:00 on 08/09/2014, Hard drive number 199.

All sensors sent off now.

comment:27 Changed 4 years ago by tec

Archiving
Beginning archiving of data.

comment:28 Changed 4 years ago by tec

Archiving
Finished archiving as of 13-01-2015 08:00

Last edited 4 years ago by tec (previous) (diff)

comment:29 Changed 4 years ago by tec

  • Description modified (diff)

comment:30 Changed 4 years ago by tec

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.